# Hearing More with Less: Multi-Modal Retrieval-and-Selection Augmented Conversational LLM-Based ASR

### v2

@@ Introduction
Automatic speech recognition (ASR) aims to convert human speech content into corresponding text. With the development of applications such as speech dialogue systems and meeting transcription, conversational ASR is becoming increasingly crucial. Typical ASR scenarios involve close-talk single speaker speech, mainly from telephone or audiobook recordings. However, conversational speech reflects the complexity of human communication, including diverse speaking styles such as specific speaker pronunciations and vocabulary preferences, paralinguistic phenomena like fillers, stutters, and repairs, as well as a high degree of contextual relevance. Previous studies indicate that incorporating speech and text modality context from preceding utterances can significantly improve conversational ASR performance (wei2024conversational; cui2023towards; shon2024generative; gong2023longfnt; hou2022bring; zhang2025mamba).
@@
Large language models (LLMs) have demonstrated exceptional long-context understanding and reasoning abilities (touvron2023llama1; touvron2023llama2; bai2023qwen), making them promising components for conversational ASR. Recent studies combining LLMs with conversational ASR involve extending LLM-based ASR (LLM-ASR) models (geng2024unveiling; bai2024seed; mu2025hdmole) by additionally inputting two types of context: a fixed number of preceding utterances (peng2025bi; li2025seewo) and the entire conversation history (bai2024seed; ding2025kimi; xu2025qwen2). The former methods assume that the most relevant historical context to the current utterance appears in the preceding few utterances, overlooking the fact that these utterances contain a large number of fillers and other semantically insignificant contexts, which can severely affect the ASR of the current utterance. Figure 1 shows that the position of the most relevant historical context for the current utterance is not fixed; it may be located in earlier conversation history, beyond a fixed number of preceding utterances. The latter methods use the entire conversation history as context. Although this provides richer context, it inevitably leads to redundant information that interferes with the ASR and incurs significant computational costs. These limitations highlight the need for a more efficient and contextually effective strategy for leveraging conversational history. Consequently, the research question can be summarized as follows: How can we retrieve and select the most relevant historical context for the current utterance to augment conversational LLM-ASR performance.
@@
Retrieval Augmented Generation (RAG) provides a potential solution to conversational LLM-ASR because it can enhance accurate, up-to-date, verifiable, and domain-specific text generation by integrating large-scale external knowledge retrieval systems with LLMs (arslan2024survey; fan2024survey; wu2024retrieval; mei2025survey; abootorabi2025ask; ni2025towards). However, RAG exhibits limited adaptability in conversational ASR. RAG focuses on generating new content based on retrieved text, whereas ASR aims to map speech to text. Thus, the two objectives are fundamentally different. Furthermore, the vast amount of content retrieved by RAG can lead to information overload in ASR, drowning out the speech that needs to be recognized. Nevertheless, the RAG design philosophy can be applied to retrieve and select the best historical context from conversational speech to assist in recognizing the current utterance.
@@
This paper proposes a multi-modal retrieval-and-selection method named MARS for enhancing conversational LLM-ASR. MARS retrieves and selects a historical context of comparable length to the current utterance, removing redundancy while retaining the most relevant information from the whole history. Specifically, multi-modal retrieval obtains a set of candidate historical contexts, each exhibiting high acoustic or textual similarity to the current utterance. Subsequently, multi-modal selection calculates the acoustic and textual similarities for each retrieved candidate historical context, and our proposed near-ideal ranking method considers both similarities and selects the best historical context. By simultaneously inputting the hypothesis from the best historical context and the current utterance speech embedding with its hypothesis into LLM-ASR, optimal ASR performance can be achieved. Experiments on the Interspeech 2025 Multilingual Conversational Speech Language Model (MLC-SLM) Challenge dataset (mu2025summary) validate the effectiveness of MARS. It achieves a significantly lower Mixed Error Rate (MER) using only 1.5K hours of training data, outperforming the top-ranking method in the challenge, which employs a specially designed LLM-ASR trained on 179K hours of data. It is worth noting that MARS sets a new state-of-the-art on the MLC-SLM dataset.
@@
Previous LLM-ASR research primarily focused on isolated utterances due to real-world conversational ASR data scarcity. The release of the MLC-SLM dataset has gradually invigorated research in conversational LLM-ASR. The TEA-ASLP system enhances Ideal-LLM (xue2024ideal) with language identification and multilingual Mixture-of-Experts (MoE), achieving optimal performance after pretraining on 179K hours of multilingual data and fine-tuning on 1.5K hours of MLC-SLM data. However, their attempt to incorporate historical and future context during the fine-tuning phase yielded no benefits (xue2025teaaslp). The Seewo system explores the upper bound of potential benefits from historical context by utilizing the ground-truth content of the two preceding utterances (li2025seewo). peng2025bi investigates the impact of using both historical and future context on conversational LLM-ASR. It proposes a character-level context masking strategy during training, where portions of the context are randomly removed to enhance robustness and better simulate potential faulty transcriptions that may occur during inference.
@@
RAG leverages external knowledge to make LLMs generate more accurate and contextually relevant content, with widespread applications in the speech modality. SEAL (sun2025seal) explores using a shared embedding space for speech-to-text retrieval, focusing on evaluating the quality of learned embeddings through proxy tasks. WavRAG (chen2025wavrag) leverages the speech encoding abilities of Qwen-Audio (chu2023qwen) to generate semantically rich speech embeddings for retrieval. feng2025enhancing proposes an RAG framework for speech-to-speech dialogue models, which can retrieve textual knowledge with input speech. pusateri2025retrieval uses a RAG-like technique to correct ASR entity name errors by querying a vector database with error-prone ASR hypotheses and feeding the retrieved entities and hypotheses to an adapted LLM for error correction. SpeechT-RAG (zhang2025speecht) improves depression detection by using speech timing features as the basis for RAG.
@@
The objective of MARS is to determine the best historical context for the current utterance through the multi-modal retrieval-and-selection method, thereby improving ASR accuracy. The overall framework of MARS is illustrated in Figure 2. The database is constructed using fully fine-tuned Whisper. It stores a triplet for each utterance in conversational speech, consisting of the utterance’s ID, speech embedding, and hypothesis. In MARS, speech embeddings and hypotheses serve as speech and text queries, respectively. The multi-modal retrieval module uses these queries to retrieve the Top-K most similar historical contexts from the database. Even after retrieval, multiple historical contexts still cause ASR prediction confusion due to information redundancy, and excessively long historical contexts will consume significant computational costs. The multi-modal selection module determines the best historical context from the retrieved contexts. Then, a language-aware prompt, the best retrieved historical context, the current utterance speech embedding, and its hypothesis input to the LLM for jointly training the projector and low-rank adaptation (LoRA) parameters (hu2022lora) to predict the transcription.
@@
In conversational ASR, both speech and text modalities carry crucial contextual information. Relying on a single modality alone cannot comprehensively retrieve historical contexts similar to the current utterance in a conversation (wei2022leveraging). In contrast, multi-modal retrieval measures the similarity of historical contexts from both speech and text perspectives, providing historical contexts for the current utterance from both pronunciation and semantic aspects. Historical contexts retrieved by speech modality can reduce ASR errors caused by pronunciation variations, while historical contexts retrieved by text modality can reduce ASR errors caused by word ambiguities. Figure 3 illustrates the detailed pipeline of the multi-modal retrieval module.
@@
In speech modality retrieval, we use the Dynamic Time Warping (DTW) to calculate the frame-level acoustic similarity between the current and historical speech embeddings. DTW calculates the matching path between two speech embeddings to determine their minimum cumulative distance. However, traditional DTW is highly computationally complex. Retrieving a large number of historical speech embeddings for the current utterance can take significant time. Therefore, we use FastDTW (salvador2007toward), which significantly reduces computational complexity while maintaining high accuracy. Additionally, we calculate utterance-level acoustic similarity by pooling the current and historical speech embeddings and computing their cosine similarity. After weighted summing the frame-level and utterance-level similarities, we obtain the speech retrieval similarities of the historical speech embeddings relative to the current utterance. We select the Top-K historical contexts with the highest speech retrieval similarities as the Top-K speech-retrieved historical contexts for the current utterance.
@@
We conduct our experiments on the MLC-SLM dataset, which originates from the recently held Interspeech 2025 Multilingual Conversational Speech Language Model Challenge [^1] https://www.nexdata.ai/competition/mlc-slm. The dataset comprises 11 languages: English, French, German, Italian, Portuguese, Spanish, Japanese, Korean, Russian, Thai, and Vietnamese. The English subset comprises approximately 500 hours of recordings from various regions, including British, American, Australian, Indian, and Philippine English. Other languages contribute around 100 hours each, resulting in a total of approximately 1500 hours of multilingual conversational speech data. Each recording consists of a natural and fluent conversation between two speakers on randomly assigned topics, recorded using devices such as iPhones in a quiet indoor environment.
@@
Implementation Details
We construct the database using the Whisper-large-v3 fully fine-tuned on the MLC-SLM dataset. For the MARS, we use the fine-tuned Whisper encoder, and the input hypotheses are pre-generated through inference by the fine-tuned Whisper model. The projector consists of two linear layers connected by a ReLU activation function. Moreover, we employ Qwen2.5-7B-Instruct (team2024qwen2) as the LLM, with LoRA configured with a rank of 64, an alpha value of 256, and a dropout rate of 0.05. Seven modules in each LLM layer, including q_proj, k_proj, v_proj, o_proj, up_proj, gate_proj, and down_proj, are subject to the LoRA. In speech modality retrieval, the frame-level and utterance-level similarities are summed with weights of 0.5 each. In text modality retrieval, we use the Qwen3-Embedding-0.6B (zhang2025qwen3) to calculate the similarities. All language-aware prompts convey the same meaning: “Please transcribe the speech into text”. The specific language prompt chosen corresponds to the language identification of the utterance. The multi-modal retrieval module generates the Top-3 speech-retrieved and text-retrieved historical contexts separately. When training MARS, we randomly mask the best retrieved historical context with a 50% probability. The MARS is trained on 6 NVIDIA A800 GPUs, with a maximum batch size accommodating 30 seconds of speech and a gradient accumulation of 6. The Adam optimizer is used with a warm-up scheduler that adjusts the learning rate, peaking at 0.0001 after 200 steps. Each model is trained for 3 epochs, and all checkpoints are averaged for the final inference. During inference, the LLM generates transcriptions without employing sampling methods, with the beam size, temperature, repetition penalty, and length penalty all set to 1.
@@
Following common evaluation standards for multilingual ASR, for languages with character-based writing systems and no clear word boundaries—including Japanese, Korean, and Thai—we use Character Error Rate (CER) to measure ASR performance. For all other languages, we use Word Error Rate (WER). Additionally, we use the Mixed Error Rate (MER) to measure the average ASR error rate across 11 languages. To ensure a fair comparison with the solutions from the MLC-SLM challenge, we use the MeetEval toolkit (von2023meeteval) to calculate all ASR error rates.
@@
Table 1 presents the WER and CER results for each language, as well as the average MER results across all languages for MARS and other comparative methods, including: 1) Vanilla Whisper-large-v3 (radford2023robust), which demonstrates excellent ASR performance across a wide range of languages; 2) Fully Fine-tuned Whisper-large-v3, which fine-tuned on the MLC-SLM training dataset; 3) Qwen2-Audio (chu2024qwen2), an speech LLM adaptable to various speech tasks and showing strong ASR performance; and 4) TEA-ASLP (xue2025teaaslp), an LLM-ASR model trained on 179K hours of multilingual ASR data including the MLC-SLM training dataset, achieved the state-of-the-art performance in the MLC-SLM test dataset. The vanilla Whisper-large-v3 demonstrates good ASR performance on the MLC-SLM dev and test datasets, and fine-tuning further improves its performance. Qwen2-Audio performs relatively well on the English subsets, but its performance on other languages is inferior due to insufficient multilingual training data. TEA-ASLP demonstrates excellent ASR performance across all languages after large-scale training. With the support of a multilingual LLM, it outperforms the fine-tuned Whisper model. MARS, using only the 1.5K hours MLC-SLM training dataset, outperforms TEA-ASLP in the majority of languages by effectively leveraging relevant historical context in conversational speech. MARS demonstrates the significant potential of retrieving and selecting suitable historical context to augment conversational ASR, highlighting remarkable data utilization that achieves high accuracy with significantly less training data.
Table 2 illustrates the comparison results of MARS with other methods that augment conversational LLM-ASR using context. Bi-context (peng2025bi) reports the results of using two preceding contexts and one future context, while Seewo (li2025seewo) reports the results of using two preceding contexts. They also evaluate ground-truth transcriptions as context to explore the upper bound of their methods. Even with ground-truth transcription as context, the benefits are limited, indicating that the immediate preceding utterances still contain irrelevant and redundant information. Furthermore, the relative gains they achieved from utilizing context are inferior to those of MARS, even when using ground-truth transcriptions. The results further underscore the necessity of MARS in retrieving and selecting the best historical context.

### v3

@@ Introduction
Automatic speech recognition (ASR) aims to convert human speech content into corresponding text. With the development of applications such as speech dialogue systems and meeting transcription, conversational ASR is becoming increasingly crucial. Typical ASR scenarios involve close-talk single speaker speech, mainly from telephone or audiobook recordings. However, conversational speech reflects the complexity of human communication, including diverse speaking styles such as specific speaker pronunciations and vocabulary preferences, paralinguistic phenomena like fillers, stutters, and repairs, as well as a high degree of contextual relevance. Previous studies indicate that incorporating speech and text modality context from preceding utterances can significantly improve conversational ASR performance (Wei et al. 2024; Cui et al. 2023; Shon et al. 2024; Gong et al. 2023; Hou et al. 2022; Zhang et al. 2025b).
@@
Large language models (LLMs) have demonstrated exceptional long-context understanding and reasoning abilities (Touvron et al. 2023a, b; Bai et al. 2023), making them promising components for conversational ASR. Recent studies combining LLMs with conversational ASR involve extending LLM-based ASR (LLM-ASR) models (Geng et al. 2024; Bai et al. 2024; Mu et al. 2025b) by additionally inputting two types of context: a fixed number of preceding utterances (Peng et al. 2025; Li et al. 2025) and the entire conversation history (Bai et al. 2024; Ding et al. 2025; Xu et al. 2025). The former methods assume that the most relevant historical context to the current utterance appears in the preceding few utterances, overlooking the fact that these utterances contain a large number of fillers and other semantically insignificant contexts, which can severely affect the ASR of the current utterance. Figure 1 shows that the position of the most relevant historical context for the current utterance is not fixed; it may be located in earlier conversation history, beyond a fixed number of preceding utterances. The latter methods use the entire conversation history as context. Although this provides richer context, it inevitably leads to redundant information that interferes with the ASR and incurs significant computational costs. These limitations highlight the need for a more efficient and contextually effective strategy for leveraging conversational history. Consequently, the research question can be summarized as follows: How can we retrieve and select the most relevant historical context for the current utterance to augment conversational LLM-ASR performance.
@@
Retrieval Augmented Generation (RAG) provides a potential solution to conversational LLM-ASR because it can enhance accurate, up-to-date, verifiable, and domain-specific text generation by integrating large-scale external knowledge retrieval systems with LLMs (Arslan et al. 2024; Fan et al. 2024; Wu et al. 2024; Mei et al. 2025; Abootorabi et al. 2025; Ni et al. 2025). However, RAG exhibits limited adaptability in conversational ASR. RAG focuses on generating new content based on retrieved text, whereas ASR aims to map speech to text. Thus, the two objectives are fundamentally different. Furthermore, the vast amount of content retrieved by RAG can lead to information overload in ASR, drowning out the speech that needs to be recognized. Nevertheless, the RAG design philosophy can be applied to retrieve and select the best historical context from conversational speech to assist in recognizing the current utterance.
@@
This paper proposes a multi-modal retrieval-and-selection method named MARS for enhancing conversational LLM-ASR. MARS retrieves and selects a historical context of comparable length to the current utterance, removing redundancy while retaining the most relevant information from the whole history. Specifically, multi-modal retrieval obtains a set of candidate historical contexts, each exhibiting high acoustic or textual similarity to the current utterance. Subsequently, multi-modal selection calculates the acoustic and textual similarities for each retrieved candidate historical context, and our proposed near-ideal ranking method considers both similarities and selects the best historical context. By simultaneously inputting the hypothesis from the best historical context and the current utterance speech embedding with its hypothesis into LLM-ASR, optimal ASR performance can be achieved. Experiments on the Interspeech 2025 Multilingual Conversational Speech Language Model (MLC-SLM) Challenge dataset (Mu et al. 2025a) validate the effectiveness of MARS. It achieves a significantly lower Mixed Error Rate (MER) using only 1.5K hours of training data, outperforming the top-ranking method in the challenge, which employs a specially designed LLM-ASR trained on 179K hours of data. It is worth noting that MARS sets a new state-of-the-art on the MLC-SLM dataset.
@@
Previous LLM-ASR research primarily focused on isolated utterances due to real-world conversational ASR data scarcity. The release of the MLC-SLM dataset has gradually invigorated research in conversational LLM-ASR. The TEA-ASLP system enhances Ideal-LLM (Xue et al. 2024) with language identification and multilingual Mixture-of-Experts (MoE), achieving optimal performance after pretraining on 179K hours of multilingual data and fine-tuning on 1.5K hours of MLC-SLM data. However, their attempt to incorporate historical and future context during the fine-tuning phase yielded no benefits (Xue et al. 2025). The Seewo system explores the upper bound of potential benefits from historical context by utilizing the ground-truth content of the two preceding utterances (Li et al. 2025). Peng et al. investigates the impact of using both historical and future context on conversational LLM-ASR. It proposes a character-level context masking strategy during training, where portions of the context are randomly removed to enhance robustness and better simulate potential faulty transcriptions that may occur during inference.
@@
RAG leverages external knowledge to make LLMs generate more accurate and contextually relevant content, with widespread applications in the speech modality. SEAL (Sun et al. 2025) explores using a shared embedding space for speech-to-text retrieval, focusing on evaluating the quality of learned embeddings through proxy tasks. WavRAG (Chen et al. 2025) leverages the speech encoding abilities of Qwen-Audio (Chu et al. 2023) to generate semantically rich speech embeddings for retrieval. Feng et al. proposes an RAG framework for speech-to-speech dialogue models, which can retrieve textual knowledge with input speech. Pusateri et al. uses a RAG-like technique to correct ASR entity name errors by querying a vector database with error-prone ASR hypotheses and feeding the retrieved entities and hypotheses to an adapted LLM for error correction. SpeechT-RAG (Zhang et al. 2025a) improves depression detection by using speech timing features as the basis for RAG.
@@
The objective of MARS is to determine the best historical context for the current utterance through the multi-modal retrieval-and-selection method, thereby improving ASR accuracy. The overall framework of MARS is illustrated in Figure 2. The database is constructed using fully fine-tuned Whisper. It stores a triplet for each utterance in conversational speech, consisting of the utterance’s ID, speech embedding, and hypothesis. In MARS, speech embeddings and hypotheses serve as speech and text queries, respectively. The multi-modal retrieval module uses these queries to retrieve the Top-K most similar historical contexts from the database. Even after retrieval, multiple historical contexts still cause ASR prediction confusion due to information redundancy, and excessively long historical contexts will consume significant computational costs. The multi-modal selection module determines the best historical context from the retrieved contexts. Then, a language-aware prompt, the best retrieved historical context, the current utterance speech embedding, and its hypothesis input to the LLM for jointly training the projector and low-rank adaptation (LoRA) parameters (Hu et al. 2022) to predict the transcription.
@@
In conversational ASR, both speech and text modalities carry crucial contextual information. Relying on a single modality alone cannot comprehensively retrieve historical contexts similar to the current utterance in a conversation (Wei et al. 2022). In contrast, multi-modal retrieval measures the similarity of historical contexts from both speech and text perspectives, providing historical contexts for the current utterance from both pronunciation and semantic aspects. Historical contexts retrieved by speech modality can reduce ASR errors caused by pronunciation variations, while historical contexts retrieved by text modality can reduce ASR errors caused by word ambiguities. Figure 3 illustrates the detailed pipeline of the multi-modal retrieval module.
@@
In speech modality retrieval, we use the Dynamic Time Warping (DTW) to calculate the frame-level acoustic similarity between the current and historical speech embeddings. DTW calculates the matching path between two speech embeddings to determine their minimum cumulative distance. However, traditional DTW is highly computationally complex. Retrieving a large number of historical speech embeddings for the current utterance can take significant time. Therefore, we use FastDTW (Salvador and Chan 2007), which significantly reduces computational complexity while maintaining high accuracy. Additionally, we calculate utterance-level acoustic similarity by pooling the current and historical speech embeddings and computing their cosine similarity. After weighted summing the frame-level and utterance-level similarities, we obtain the speech retrieval similarities of the historical speech embeddings relative to the current utterance. We select the Top-K historical contexts with the highest speech retrieval similarities as the Top-K speech-retrieved historical contexts for the current utterance.
@@
We conduct our experiments on the MLC-SLM dataset, which originates from the recently held Interspeech 2025 Multilingual Conversational Speech Language Model Challenge (Mu et al. 2025a). The dataset comprises 11 languages: English, French, German, Italian, Portuguese, Spanish, Japanese, Korean, Russian, Thai, and Vietnamese. The English subset comprises approximately 500 hours of recordings from various regions, including British, American, Australian, Indian, and Philippine English. Other languages contribute around 100 hours each, resulting in a total of approximately 1500 hours of multilingual conversational speech data. Each recording consists of a multi-turn, natural, and fluent conversational speech of around 20 minutes between two speakers on a randomly assigned topic, including celebrities, dreams, education, emotion, fashion, food, games, the Internet, movies, shopping, travel, etc, recorded using devices such as iPhones in a quiet indoor environment.
Implementation Details
We construct the database using the Whisper-large-v3 fully fine-tuned on the MLC-SLM dataset. For the MARS, we use the fine-tuned Whisper encoder, and the input hypotheses are pre-generated through inference by the fine-tuned Whisper model. The projector consists of two linear layers connected by a ReLU activation function. Moreover, we employ Qwen2.5-7B-Instruct (Yang et al. 2024) as the LLM, with LoRA configured with a rank of 64, an alpha value of 256, and a dropout rate of 0.05. Seven modules in each LLM layer, including q_proj, k_proj, v_proj, o_proj, up_proj, gate_proj, and down_proj, are subject to the LoRA. In speech modality retrieval, the frame-level and utterance-level similarities are summed with weights of 0.5 each. In text modality retrieval, we use the Qwen3-Embedding-0.6B (Zhang et al. 2025c) to calculate the similarities.
@@
All language-aware prompts convey the same meaning: “Please transcribe the speech into text”. The specific language prompt chosen corresponds to the language identification of the utterance. The multi-modal retrieval module generates the Top-3 speech-retrieved and text-retrieved historical contexts separately. When training MARS, we randomly mask the best retrieved historical context with a 50% probability. The MARS is trained on 6 NVIDIA A800 GPUs, with a maximum batch size accommodating 30 seconds of speech and a gradient accumulation of 6. The Adam optimizer is used with a warm-up scheduler that adjusts the learning rate, peaking at 0.0001 after 200 steps. Each model is trained for 3 epochs, and all checkpoints are averaged for the final inference. During inference, the LLM generates transcriptions without employing sampling methods, with the beam size, temperature, repetition penalty, and length penalty all set to 1.
@@
Following common evaluation standards for multilingual ASR, for languages with character-based writing systems and no clear word boundaries—including Japanese, Korean, and Thai—we use Character Error Rate (CER) to measure ASR performance. For all other languages, we use Word Error Rate (WER). Additionally, we use the Mixed Error Rate (MER) to measure the average ASR error rate across 11 languages. To ensure a fair comparison with the solutions from the MLC-SLM challenge, we use the MeetEval toolkit (von Neumann et al. 2023) to calculate all ASR error rates.
@@
Table 1 presents the WER and CER results for each language, as well as the average MER results across all languages for MARS and other comparative methods, including: 1) Vanilla Whisper-large-v3 (Radford et al. 2023), which demonstrates excellent ASR performance across a wide range of languages; 2) Fully Fine-tuned Whisper-large-v3, which fine-tuned on the MLC-SLM training dataset; 3) Qwen2-Audio (Chu et al. 2024), an speech LLM adaptable to various speech tasks and showing strong ASR performance; and 4) TEA-ASLP (Xue et al. 2025), an LLM-ASR model trained on 179K hours of multilingual ASR data including the MLC-SLM training dataset, achieved the state-of-the-art performance in the MLC-SLM test dataset. The vanilla Whisper-large-v3 demonstrates good ASR performance on the MLC-SLM dev and test datasets, and fine-tuning further improves its performance. Qwen2-Audio performs relatively well on the English subsets, but its performance on other languages is inferior due to insufficient multilingual training data. TEA-ASLP demonstrates excellent ASR performance across all languages after large-scale training. With the support of a multilingual LLM, it outperforms the fine-tuned Whisper model. MARS, using only the 1.5K hours MLC-SLM training dataset, outperforms TEA-ASLP in the majority of languages by effectively leveraging relevant historical context in conversational speech. MARS demonstrates the significant potential of retrieving and selecting suitable historical context to augment conversational ASR, highlighting remarkable data utilization that achieves high accuracy with significantly less training data.
Table 2 illustrates the comparison results of MARS with other methods that augment conversational LLM-ASR using context. Bi-context (Peng et al. 2025) reports the results of using two preceding contexts and one future context, while Seewo (Li et al. 2025) reports the results of using two preceding contexts. They also evaluate ground-truth transcriptions as context to explore the upper bound of their methods. Even with ground-truth transcription as context, the benefits are limited, indicating that the immediate preceding utterances still contain irrelevant and redundant information. Furthermore, the relative gains they achieved from utilizing context are inferior to those of MARS, even when using ground-truth transcriptions. The results further underscore the necessity of MARS in retrieving and selecting the best historical context.